OpenStudy (kaitlyn_nicole):

Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)).

OpenStudy (anonymous):

try it, i will check

OpenStudy (vuriffy):

You must have some knowledge of this topic, so at least attempt the problem. Satellite and I may help you in anyway possible to ensure you understand the concept.

OpenStudy (boldjon):

f(x) = -4x + 7 g(x) = 10x - 6 To get f(g(x)), simply replace the x in f(x) with g(x): f(g(x)) = -4g(x) + 7 = -4(10x-6) + 7 = -40x + 24 + 7

OpenStudy (boldjon):

do the last part and that's yo answer :3

OpenStudy (vuriffy):

Boldjon, please give them a chance to attempt the question, but thank you for your help.

OpenStudy (kaitlyn_nicole):

okay, i was thinking -4(10-6)+7

OpenStudy (boldjon):

hm... yeah that was close, but the answer es = -40x + 31 HARIBAAAA ur welcome gtg :3

OpenStudy (kaitlyn_nicole):

so then you distribute, and get -40x+31

OpenStudy (anonymous):

yeah but you need an \(x\) next to the \(10\)

OpenStudy (boldjon):

yes ur right :3

OpenStudy (kaitlyn_nicole):

ooh ok

OpenStudy (anonymous):

not \[ -4(10-6)+7\] but rather \[ -4(10x-6)+7\]

OpenStudy (kaitlyn_nicole):

lright! so what happens when they give you an extra number; for example: Let f(x) = x − 2 and g(x) = x^2 − 7x − 9. Find f(g(−1)).

OpenStudy (anonymous):

it is actually easier, no algebra, just arithmetic

OpenStudy (kaitlyn_nicole):

when you get it would you just plug in the x to get a number?

OpenStudy (anonymous):

first compute \(g(-1)\) what do you get?

OpenStudy (kaitlyn_nicole):

ummm, 1+7-9? is that what you meant??

OpenStudy (anonymous):

yes, for a grand total of?

OpenStudy (kaitlyn_nicole):

-1? oh what lol

OpenStudy (kaitlyn_nicole):

it equals the same?

OpenStudy (anonymous):

yes -1

OpenStudy (anonymous):

now compute \(f(-1)\)

OpenStudy (kaitlyn_nicole):

-3?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so \[f(g(-1))=f(-1)=-3\] that is your answer

OpenStudy (kaitlyn_nicole):

o wow! ill try one now

OpenStudy (anonymous):

k good

OpenStudy (kaitlyn_nicole):

Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). g(x)= 1 f(x)=10 so, f(g(2))=g(1)=10?

OpenStudy (anonymous):

i think your \(g(2)\) is wrong, but you have the right idea

OpenStudy (kaitlyn_nicole):

oh, oops!

OpenStudy (kaitlyn_nicole):

i have one more question and that is it for functions!

OpenStudy (anonymous):

did you get the final answer to this one?

OpenStudy (kaitlyn_nicole):

10

OpenStudy (kaitlyn_nicole):

Let f(x) = 5x + 12. Find f−1(x). we did one of these right?

OpenStudy (anonymous):

yes but we did it with numbers now it is with variables change x and y, solve \[x=5y+12\]for \(y\) in two steps

OpenStudy (boldjon):

By f-1(x) you mean the inverse I assume. To do that just swap the values of your function with y<->x. y=5x+12 becomes x=5y+12 and now solve for y.

OpenStudy (boldjon):

if you had something like 1/5x - 12/5, that would be the closest

OpenStudy (kaitlyn_nicole):

ok so -12 + x=5y?

OpenStudy (anonymous):

yes that is step one

OpenStudy (kaitlyn_nicole):

both of you explaining is so difficult lol

OpenStudy (kaitlyn_nicole):

ok! so would i divide by 5?

OpenStudy (anonymous):

eys

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

by divide, just write it

OpenStudy (kaitlyn_nicole):

so \[\frac{ x-12 }{ 5 }\]

OpenStudy (anonymous):

bingo

OpenStudy (kaitlyn_nicole):

wooo thank you! I got a 90

OpenStudy (anonymous):

yay (i guess)

OpenStudy (anonymous):

100 would have been better

OpenStudy (kaitlyn_nicole):

yeah inknow! i got the 2nd to last one wrong

OpenStudy (kaitlyn_nicole):

that my g(2) was messed up for something?

OpenStudy (anonymous):

prolly

OpenStudy (kaitlyn_nicole):

but i got the curriculum, so thank you for taking your time and explaining to it me